# Selective Variable Convolution Meets Dynamic Content Guided Attention for Infrared Small Target Detection

### v1

@@ Abstract
Infrared Small Target Detection (IRSTD) system aims to identify small targets in complex backgrounds. Due to the convolution operation in Convolutional Neural Networks (CNNs), applying traditional CNNs to IRSTD presents challenges, since the feature extraction of small targets is often insufficient, resulting in the loss of critical features. To address these issues, we propose a dynamic content guided attention multiscale feature aggregation network (DCGANet), which adheres to the attention principle of ’coarse-to-fine’ and achieves high detection accuracy. First, we propose a selective variable convolution (SVC) module that integrates the benefits of standard convolution, irregular deformable convolution, and multi-rate dilated convolution. This module is designed to expand the receptive field and enhance non-local features, thereby effectively improving the discrimination of targets from backgrounds. Second, the core component of DCGANet is a two-stage content guided attention module. This module employs two-stage attention mechanism to initially direct the network’s focus to salient regions within the feature maps and subsequently determine whether these regions correspond to targets or background interference. By retaining the most significant responses, this mechanism effectively suppresses false alarms. Additionally, we propose adaptive dynamic feature fusion (ADFF) module to substitute for static feature cascading. This dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, thereby enhancing its ability to discriminate true targets from false alarms. DCGANet has achieved new benchmarks across multiple datasets.
@@
Infrared small target, selective variable convolution, dynamic content guided attention, adaptive dynamic feature fusion.
@@ I Introduction
Infrared imaging technology, a crucial component of infrared monitoring systems, excels at capturing thermal signatures of targets and demonstrates robust performance in low-light and foggy environments. In military [1], ocean rescue [2], and environmental monitoring [3] applications, targets of interest at long imaging distances often appear as small spots [3]. Consequently, infrared small target detection (IRSTD) is essential for extracting valuable information from complex backgrounds.
Figure 1: The complex background faced by image target detection, the red and green bounding boxes are zoom in area: (a)Dense target of different types such as pedestrians and cars, (b)Target occlusion by other objects such as trees, (c)Strong lighting conditions of sunlight, and (d) Dim lighting conditions during nights.
@@
Addressing these challenges is essential for achieving robust and precise detection in infrared warning systems. It is imperative to suppress background clutter by enhancing the extraction of inherent target features and integrating contextual analysis, which is pivotal to improving IRSTD performance.
Traditional IRSTD methodologies can be categorized into three primary approaches. They are Background Suppression-based (BS-based), Human Visual System-based (HVS-based), and Low Rank Matrix-based (LRM-based) methods. The BS-based methods [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] utilize specialized filters to extract small targets. The HVS-based methods [16, 17, 18, 19, 20, 21, 22, 23] identify small targets by exploiting local information features. The LRM-based methods [24, 25, 26, 27] detect small targets through the decomposition and subsequent reconstruction of infrared images into sparse matrices. However, these traditional methods are limited in their ability to learn from datasets or dynamically adjust parameters in response to changing environmental conditions, leading to suboptimal performance in complex scenarios.
Recent advances in deep learning have yielded substantial improvements for IRSTD. Previous research in deep learning methods emphasized weak feature expressions, such as gradient edge [28, 29, 30, 31], contextual information [32, 33, 34, 35, 36], and local contrast [37, 38, 39]. Despite their progress, fundamental challenges in IRSTD remain unaddressed. Specifically, the current limitations are primarily evident in three aspects:
@@
    Static Attention for Interference Suppression: Traditional convolution and pooling strategies often struggle to effectively separate background and target features, while the saliency of infrared targets may be distributed in specific channels (such as thermal radiation intensity, gradient changes, etc.). However, existing attention mechanisms (such as SE [10], CBAM [34]) model channel relationships statically or at a single scale, causing target information to be masked by background signals, reducing detection sensitivity.
@@
Facing the challenge of IRSTD, existing convolutional neural networks often struggle to effectively address the feature extraction of infrared small targets. To overcome these challenges, there is an imperative need for a novel framework, which is derived from three pivotal observations: (1) Infrared small targets require hierarchical dynamic receptive fields to jointly capture edge sharpness and thermal radiation context; (2) Feature fusion should be selectively guided by high-level semantics to suppress noise propagation; (3) Attention mechanisms must incorporate infrared physical priors and consider input channels according to wavelength sensitivity, while adapting to environmental dynamics. Therefore, our framework should not only effectively extract subtle features of infrared small targets through advanced feature extraction techniques but also dynamically optimize multi-stage feature fusion processes. This framework requires a deep understanding of the background environment, rather than simply relying on local image information near the target.
To address the aforementioned challenges, we propose DGCANet, which consists of three core innovations. First, to improve the recognition of real targets, suppress background noise, and discriminate between them and false alarms, we proposed selective variable convolution (SVC), which consists of multi-branch adaptive kernels. The SVC resolving the local-global trade-off, which can accurately extract features of weak targets and significantly reduce background noise. Second, a dynamic content guided attention(DCGA) mechanism was proposed. The DGCA first performs local feature extraction and expanded receptive field extraction using parallel ordinary convolution and dilated convolution. To generate multi-scale coarse-grained attention maps, through channel grouping refinement strategy, the input feature channels are grouped according to physical meanings (such as thermal radiation intensity and gradient texture), and captures insensitive features of weak targets through fine attention allocation. Last, we proposed Adaptive Dynamic Feature Fusion (ADFF) to simulate the attention regulation mechanism of the human vision systems(HVS). The integration of ADFF with a wide receptive field significantly improves the distinguish between real targets and false positives.
@@
    We propose DCGANet, a novel framework that unifies feature learning, dynamic content guided attention, and feature fusion in a hierarchical dynamic pipeline. This framework bridges the gap between rigid CNN designs and the dynamic, sparse infrared target, achieving joint optimization of accuracy, robustness, and interpretability.
@@ II-A1 Model-Driven Infrared Small Target Detection
Model-driven methods formalize IRSTD as anomaly recognition against backgrounds, leveraging three computational paradigms: filters-based, human visual system (HVS) simulations-based, and sparse matrix reconstruction. The filter-based method [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] mainly focuses on suppressing background clutter and highlighting targets while suppressing isolated noise through filter stacking. For instance, multi-scale Top-Hat [8, 9] filtering algorithm, which combines multi-scale transformation, can improve detection performance. However, when the infrared image contains complex backgrounds, the algorithm will exhibit a lower signal-to-noise ratio (SNR). To further improve the performance, the Human Visual System-Based (HVS-Based) methods [16, 17, 18, 19, 20, 21, 22, 23, 41] utilize the principle of human visual contrast to suppress background brightness and enhance local information of infrared small targets, such as LCM [16], and TLLCM [41]. However, HVS-based methods still struggle to maintain detection stability when faced with high interference from bright edges or points. From another perspective, sparse matrix reconstruction treats small targets as sparse matrices and backgrounds as low-rank matrices [24, 25, 26, 27, 42], detecting small targets by decomposition low-rank matrices. For instance, the Infrared Patch Image (IPI) [42] model divides infrared images into blocks and recombines them into low-rank sparse matrices. Meanwhile, the reweighted infrared patch tensor (RIPT) [24] was proposed to construct low-rank sparse matrices more stably and faster through the alternating direction method. Overall, although these three categories of methods can achieve good detection results for certain images, their performance dramatically decreases when facing complex backgrounds.
@@ II-A2 Data-Driven Infrared Small Target Detection
Although deep learning methods have made substantial progress, they still face challenges such as over-parameterized architectures and inflexible receptive fields [43], impending comprehensive capture of the details of small targets. Excessive reliance on local information [31] also limits their ability to model global contextual features, leading to high false alarm rates in noisy scenarios. DCGANet distinguishes itself through two core innovations:
@@
    HVS-mimetic dynamic attention: The DCGA module implements ’coase-to-fine’ saliency mapping, mirroring human visual system search patterns through channel-specific attention refinement.
@@ II-B Vision Attention Mechanisms on IRSTD
Despite these advancements, existing methods still lack robust global feature understanding in IRSTD. In contrast, dynamic content guided attention (DCGA) implements a coarse-to-fine attention cascade: it first generates a rough spatial attention map, then refines it into a fine-grained map by integrating spatial and channel attention. This process refines initial spatial attention maps (SIM) for each channel to focus on the most representative features, enhancing the recognition of suspected small targets. In feature fusion, DCGA adopts a shallow-deep feature fusion strategy to improve information flow between network layers, preserving features and enabling effective gradient backpropagation. This dual optimization not only boosts small target detection accuracy but also strengthens false alarm suppression capabilities.
III Method
III-A Overall Architecture
@@
Figure 2: Overall architecture of the proposed DCGANet, built upon a U-Net-like backbone and integrating the key Selective Variable Convolution (SVC), Dynamic Content Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) modules.
@@
The DCGANet architecture, depicted in Fig. 2, is constructed upon a U-Net-like framework and comprises three components: selective variable convolution (SVC), dynamic content guided attention (DCGA), and adaptive dynamic feature fusion (ADFF) modules. Specifically, the encoder path consists of three cascaded layers integrating SVC and DCGA modules in sequence. The SVC replaces standard convolutions to enhance subtle target signatures while expanding receptive fields for augmented context modeling. The DCGA mechanism generates channel-specific spatial importance maps (SIMs) through input-guided attention calibration, enabling per-channel saliency emphasis critical for sub-pixel infrared targets. On the decoder path, two ADFF modules dynamically merge multi-scale encoder features, replacing the original skip-connections and achieving context-aware noise suppression through three-stage fusion.
III-B Selective Variable Convolution
@@
The limited receptive field of traditional convolutional layers in CNNs hinders their ability in IRSTD. To address this, we propose the selective variable convolution (SVC) module, a robust basic block aimed at significantly enhancing the network’s feature learning ability.
As shown in Fig. 3, the SVC combines synergy of the tri-branch features. The first branch employs standard convolution(SConv) to preserve local spatial feature extraction, retaining elementary patterns of small targets and maintaining the stability of fundamental detection performance. The second branch utilizes deformable convolution (DConv), which has been proven effective in edge delineation for irregularly shaped targets, enhancing target features across diverse scales. The third branch, multi-dilation convolution (MDConv), utilizes multi-rate dilated kernels (rates = 2, 4, 8) to extend the receptive field, thereby aggregating multi-scale contexts to suppress structured clutter. The SVC enables accurate extraction of weak target features and significant reduction of background noise.
@@
$X_{m}$ $=\operatorname{MDConv}(X_{\text{in}}).$ (3)
where $X_{\text{in}}$ is the input feature, $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of SConv, DConv, and MDConv, respectively.
The output of these branches is fused through channel-wise concatenation and a $1\times 1$ convolution, yielding consolidated representations enriched in both local precision and global discriminability.
@@
where $X_{\text{out}}$ is the output of SVC module, $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of SConv, DConv, and MDConv, respectively.
By enhancing the extraction of edges and details through a multi-convolution strategy, SVC effectively boosts recognition capability in IRSTD. The expanded receptive field and deformable feature extraction offer substantial advantages in contextual perception and noise suppression, thereby improving the clarity and reliability of feature extraction.
Figure 3: The structure of the SVC: a parallel branching architecture to perform different types of convolution operations on the input feature maps. This design enables the encoding phase to extract features and preserve details at various scales.
III-C Dynamic Content Guided Attention
While the SVC module expands the receptive field and enhances non-local feature extraction, it still struggles to identify the target when dealing with background clutter. Previous attention mechanisms, such as the static channel-wise or spatial weight separation mechanisms of SE [10], CSAM [34], face two key issues. First, there is a lack of information interaction between channel-wise and spatial attention. Second, single-channel spatial weights cannot adapt to the diverse distribution of multiple feature channels and pose a risk of information loss due to the dilution of key target features.
To address these challenges, we propose the dynamic content guided attention(DCGA) module, which uses a coarse-to-fine approach to generate self-attention maps(SIM) for specific channels, effectively simulating the ability of HVS to focus on highlighting features and suppress false alarms. The DCGA is shown in Fig. 4. The first stage includes coarse-grained attention with interaction between channel-wise and spatial attention, while the second stage includes fine-grained calibration with dynamic SIM generation, obtaining a unique SIM for each single-channel input features in a coarse-to-fine manner. We first calculate the corresponding Wc and Ws by following.
III-C1 Stage I: Coarse-Grained Attention with Joint Channel-Space Fusion
The stage include Channel attention, Spatial Attention, and joint attention.
Channel attention: A dual-bottleneck structure extracts global channel responses by aggregating spatial information through global average pooling:
$W_{c}=\operatorname{\mathcal{C}}_{1\times 1}\left(\operatorname{ReLU}\left(\operatorname{\mathcal{C}}_{1\times 1}\left(X_{GAP}^{\text{c}}\right)\right)\right)$ (5)
where $X_{GAP}^{\text{c}}$ denotes channel-wise squeezed features from global average pooling.
Spatial Attention: Spatial saliency is captured by concatenating multi-perspective features from global average pooling (GAP) and global max pooling (GMP):
$W_{s}=\operatorname{\mathcal{C}}_{7\times 7}\left(\operatorname{Concat}\left(X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}\right)\right).$ (6)
where $X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}$ represent spatial descriptors from global average pooling (GAP) and global max pooling (GMP).
Interaction Attention: we fuse $W_{c}$ and $W_{s}$ together using an element-wise addition operation following broadcast rules to obtain a rough SIM $W_{\text{coa}}$.

### v2

# Selective Variable Convolution Meets Dynamic Content-Guided Attention for Infrared Small Target Detection

@@ Abstract
Infrared Small Target Detection (IRSTD) system aims to identify small targets in complex backgrounds. Due to the convolution operation in Convolutional Neural Networks (CNNs), applying traditional CNNs to IRSTD presents challenges, since the feature extraction of small targets is often insufficient, resulting in the loss of critical features. To address these issues, we propose a dynamic content-guided attention multiscale feature aggregation network (DCGANet), which adheres to the attention principle of ’coarse-to-fine’ and achieves high detection accuracy. First, we propose a selective variable convolution (SVC) module that integrates the benefits of standard convolution, irregular deformable convolution, and multi-rate dilated convolution. This module is designed to expand the receptive field and enhance non-local features, thereby effectively improving the discrimination between targets and backgrounds. Second, the core component of DCGANet is a two-stage content-guided attention module. This module employs a two-stage attention mechanism to initially direct the network’s focus to salient regions within the feature maps and subsequently determine whether these regions correspond to targets or background interference. By retaining the most significant responses, this mechanism effectively suppresses false alarms. Additionally, we propose an Adaptive Dynamic Feature Fusion (ADFF) module to substitute for static feature cascading. This dynamic feature fusion strategy enables DCGANet to adaptively integrate contextual features, thereby enhancing its ability to discriminate true targets from false alarms. DCGANet has achieved new benchmarks across multiple datasets.
@@
Infrared small target , Selective variable convolution ,Dynamic content-guided attention ,Adaptive dynamic feature fusion
@@ I Introduction
Infrared imaging technology, a crucial component of infrared monitoring systems, excels at capturing thermal signatures of targets and demonstrates robust performance in low-light and foggy environments. In military [1], ocean rescue [2], and environmental monitoring [3, 4] applications, targets of interest at long imaging distances often appear as small spots [3]. Consequently, infrared small target detection (IRSTD) is essential for extracting valuable information from complex backgrounds.
Figure 1: Due to the long imaging distance, real targets (red boxes) lack clear texture and structure, appearing as tiny bright spots. Meanwhile, high-brightness clutter in the background, such as from clouds, buildings, or the ground (green boxes), actively mimics real targets in local morphology, size, and brightness, creating indistinguishable interference.
@@
Addressing these challenges is essential for achieving robust and precise detection in infrared warning systems. To improve IRSTD performance, it is crucial to suppress background clutter. This can be achieved by enhancing the extraction of inherent target features and integrating contextual analysis.
Traditional data-driven IRSTD methodologies can be categorized into three primary approaches. They are Background Suppression-based (BS-based), Human Visual System-based (HVS-based), and Low Rank Matrix-based (LRM-based) methods. The BS-based methods [6, 7, 8, 9] utilize specialized filters to extract small targets. The HVS-based methods [10, 11, 12] identify small targets by exploiting local information features. The LRM-based methods [13, 14] detect small targets through the decomposition and subsequent reconstruction of infrared images into sparse matrices. However, these traditional methods are limited in their ability to learn from datasets or dynamically adjust parameters in response to changing environmental conditions, leading to suboptimal performance in complex scenarios.
Recent advances in deep learning have yielded substantial improvements for data-driven IRSTD [15]. Previous research in deep learning methods emphasized weak feature expressions, such as gradient edge [16], contextual information [17, 18, 19, 20], and local contrast [21, 22]. Despite their progress, fundamental challenges in IRSTD remain unaddressed. Specifically, the current limitations are primarily evident in three aspects:
@@
    Static Attention for Interference Suppression: Traditional convolution and pooling strategies often struggle to effectively separate background and target features, while the saliency of infrared targets may be distributed in specific channels (such as thermal radiation intensity, gradient changes, etc.). However, existing attention mechanisms (such as SE [23], CBAM [19]) model channel relationships statically or at a single scale [20], causing target information to be masked by background signals, reducing detection sensitivity.
@@
Facing the challenges encountered in existing infrared small target detection, this study sets its core objective: to construct a high-performance detection framework capable of synergistically optimizing feature learning, contextual awareness, and feature fusion. Firstly, a dynamic feature extraction mechanism is introduced. This mechanism aims to resolve the inherent conflict between local receptive fields and global contextual modeling, adaptively adjusting its information aggregation scope to enhance the representational capacity of foundational features. Secondly, a multi-scale context-guided attention strategy is employed. This strategy first expands the perceptual range of local features and conducts intelligent grouping and refinement along the channel dimension based on the physical properties of infrared imaging, thereby generating multi-scale attention maps. Finally, to address missed detections and false alarms caused by improper fusion, an adaptive feature fusion method is proposed. This method dynamically evaluates and balances the contributions of features from different hierarchical levels, intelligently integrating complementary information. In summary, this research is committed to achieving breakthroughs in three critical aspects—feature extraction, information perception, and feature fusion—through the above synergistic approaches. The ultimate goal is to establish a robust infrared small target detection system capable of delivering high-precision and high robustness in complex, dynamic infrared scenarios.
To address the aforementioned challenges, we propose DCGANet, which consists of three core innovations. First, to improve the recognition of real targets, suppress background noise, and discriminate between them and false alarms, we proposed selective variable convolution (SVC), which consists of multi-branch adaptive kernels. The SVC resolves the local-global trade-off, which can accurately extract features of weak targets and significantly reduce background noise. Second, we proposed a Dynamic Content-Guided Attention (DCGA) mechanism. This mechanism simulates the "coarse-to-fine" search pattern of the human visual system. Through a two-stage process, it dynamically generates a dedicated Spatial Importance Map (SIM) for each channel of the feature map, thereby precisely enhancing target features while suppressing background interference. Last, to address the issue of improper fusion between different hierarchical features, we designed the Adaptive Dynamic Feature Fusion (ADFF) module. It replaces traditional skip connections, intelligently weighing and merging features from the encoder and decoder through a dynamic spatial selection mechanism. This effectively suppresses the contamination of deep semantic features by shallow-level noise.
@@
    We propose DCGANet, a novel framework that unifies feature learning, contextual guide attention, and feature fusion in a hierarchical dynamic pipeline. This framework bridges the gap between rigid CNN designs and the dynamic, sparse infrared target, achieving joint optimization of accuracy, robustness, and interpretability.
@@ II-A1 Model-Driven Infrared Small Target Detection
Model-driven methods formalize IRSTD as anomaly recognition against backgrounds, leveraging three computational paradigms: filters-based, human visual system (HVS) simulations-based, and sparse matrix reconstruction. The filter-based method [6, 7, 8] mainly focuses on suppressing background clutter and highlighting targets while suppressing isolated noise through filter stacking. For instance, multi-scale Top-Hat [9, 25] filtering algorithm, which combines multi-scale transformation, can improve detection performance. However, when the infrared image contains complex backgrounds, the algorithm will exhibit a lower signal-to-noise ratio (SNR). To further improve the performance, the Human Visual System-Based (HVS-Based) methods [10, 11, 12, 26] utilize the principle of human visual contrast to suppress background brightness and enhance local information of infrared small targets, such as LCM [10], and TLLCM [27]. However, HVS-based methods still struggle to maintain detection stability when faced with high interference from bright edges or points. From another perspective, sparse matrix reconstruction treats small targets as sparse matrices and backgrounds as low-rank matrices [13, 14, 28], detecting small targets by decomposition low-rank matrices. For instance, the Infrared Patch Image (IPI) [29] model divides infrared images into blocks and recombines them into low-rank sparse matrices. Meanwhile, the reweighted infrared patch tensor (RIPT) [13] was proposed to construct low-rank sparse matrices more stably and faster through the alternating direction method. Overall, although these three categories of methods can achieve good detection results for certain images, their performance dramatically decreases when facing complex backgrounds.
@@ II-A2 Data-Driven Infrared Small Target Detection
Although deep learning methods have made substantial progress, they still face challenges such as over-parameterized architectures and inflexible receptive fields [30], impending comprehensive capture of the details of small targets. Excessive reliance on local information [32, 33] also limits their ability to model global contextual features, leading to high false alarm rates in noisy scenarios. DCGANet distinguishes itself through two core innovations:
@@
    HVS-mimetic dynamic attention: The DCGA module implements ’coarse-to-fine’ saliency mapping, mirroring human visual system search patterns through channel-specific attention refinement.
@@ II-B Vision Attention Mechanisms on IRSTD
Despite these advancements, existing methods still lack robust global feature understanding in IRSTD. In contrast, Dynamic Content-Guided Attention (DCGA) implements a coarse-to-fine attention cascade: it first generates a rough spatial attention map, then refines it into a fine-grained map by integrating spatial and channel attention. This process refines initial spatial attention maps (SIM) for each channel to focus on the most representative features, enhancing the recognition of suspected small targets. In feature fusion, DCGA adopts a shallow-deep feature fusion strategy to improve information flow between network layers, preserving features and enabling effective gradient backpropagation. This dual optimization not only boosts small target detection accuracy but also strengthens false alarm suppression capabilities.
III Methodology
@@
Figure 2: The overall architecture of the proposed DCGANet. It is built upon a U-Net-like backbone and integrates the key Selective Variable Convolution (SVC), Dynamic Context-Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) modules.
Figure 3: The structure of the SVC module: a parallel branch architecture that performs different types of convolution operations on the input feature map. This design enables the encoding stage to extract features at different scales and preserve details.
@@
The overall architecture of DCGANet is illustrated in Figure 2, built upon a U-Net-like framework and composed of three core components: the Selective Variable Convolution (SVC), Dynamic Context-Guided Attention (DCGA), and Adaptive Dynamic Feature Fusion (ADFF) modules. Specifically, the encoder path consists of three cascaded layers, each sequentially integrating an SVC and a DCGA module. SVC replaces standard convolution to enhance faint target features while expanding the receptive field for stronger contextual modeling. The subsequent DCGA mechanism generates channel-specific Spatial Importance Maps (SIMs) through input-guided attention calibration, enabling per-channel saliency emphasis crucial for sub-pixel infrared targets. On the decoder path, two ADFF modules dynamically fuse multi-scale features from the encoder, replacing the original skip connections and effectively suppressing noise through a context-aware fusion strategy.
III-A Selective Variable Convolution
@@
In Convolutional Neural Networks (CNNs), the limited receptive field of traditional convolutional layers hinders their ability to capture global contextual information in Infrared Small Target Detection (IRSTD). To address this, we propose the Selective Variable Convolution (SVC) module, a powerful foundational block designed to significantly enhance the network’s feature learning capabilities.
The structure of SVC, shown in Figure 3, enhances expressive power by integrating features from three parallel branches.
  • Standard Convolution (SConv): The first branch employs a standard convolution to preserve local spatial features, maintaining the stability of basic patterns for small targets and foundational detection performance.
  • Deformable Convolution (DConv): The second branch utilizes deformable convolution, which excels at outlining the edges of irregularly shaped targets, enhancing target features at different scales.
  • Multiple Dilation Convolution (MDConv): The third branch is a multiple dilation convolution, which uses kernels with different dilation rates (rates = 2, 4, 8) to expand the receptive field, thereby aggregating multi-scale context to suppress structured clutter.
The mathematical expressions for these three branches are as follows:
@@
$X_{m}$ $=\operatorname{MDConv}(X_{\text{in}}),$ (3)
where $X_{\text{in}}$ is the input feature, and $X_{s}$, $X_{c}$, and $X_{m}$ are the outputs of the three branches, respectively.
The outputs of these branches are fused via channel-wise concatenation and a $1\times 1$ convolution, yielding an integrated representation enriched in both local precision and global differentiability.
@@
where $X_{\text{out}}$ is the final output of the SVC module. Through this multi-convolution strategy, SVC effectively improves the recognition of target edges and details, and its expanded receptive field and deformable feature extraction provide significant advantages for context-aware modeling and noise suppression.
III-B Dynamic Context-Guided Attention
Although the SVC module enhances feature extraction, it may still face difficulties when dealing with complex background clutter. Traditional attention mechanisms suffer from two key problems: first, a lack of information interaction between channel and spatial attention; and second, a single spatial weight for all channels cannot adapt to the diverse distributions across multiple feature channels, risking the dilution of key target information.
To address these challenges, we propose the Dynamic Context-Guided Attention (DCGA) module. The structure of DCGA, shown in Figure 4, employs a coarse-to-fine approach to Spatial Importance Map (SIM), effectively simulating the Human Visual System’s (HVS) ability to focus on salient features and suppress false positives.
Figure 4: The structure of the DCGA module: It employs a parallel approach to equally weight spatial and channel information. It also incorporates a channel shuffling of channel and spatial information to enhance target segmentation capabilities.
III-B1 Stage 1: Coarse-Grained Attention and Joint Channel-Spatial Fusion
This stage generates a preliminary attention map.
Channel Attention: A dual-bottleneck structure is used, aggregating spatial information via Global Average Pooling (GAP) to extract the global channel response $W_{c}$:
$W_{c}=\operatorname{\mathcal{C}}_{1\times 1}\left(\operatorname{ReLU}\left(\operatorname{\mathcal{C}}_{1\times 1}\left(X_{GAP}^{\text{c}}\right)\right)\right),$ (5)
where $X_{GAP}^{\text{c}}$ is the channel-compressed feature obtained from Global Average Pooling.
Spatial Attention: Spatial saliency $W_{s}$ is captured by concatenating multi-view features from Global Average Pooling (GAP) and Global Max Pooling (GMP):
$W_{s}=\operatorname{\mathcal{C}}_{7\times 7}\left(\operatorname{Concat}\left(X_{GAP}^{\text{s}},X_{GMP}^{\text{s}}\right)\right),$ (6)
where $X_{GAP}^{\text{s}}$ and $X_{GMP}^{\text{s}}$ represent the spatial descriptors from GAP and GMP, respectively.
Interactive Attention: Following broadcasting rules, we fuse $W_{c}$ and $W_{s}$ using element-wise addition to obtain a coarse SIM $W_{\text{coa}}$.